Abby is registering at a Web site. She must select a password
containing six numerals to be able to use the site. How many passwords are
allowed if no digit may be used more than once?
I get differennt answer from different people , Please! help me select correct answer, thank you in advance
a) 6!= 720
b) (10)!?( 10-6)!= 151.200
c) 10*9*8*7*6*5 = 151.200
d) 10+9+8+7+6+5=45
e) 10C6=10x9x8x7x6x5x4!/6x5x4x3x2x4!
=3x2x7x5=210
which one correct answer ?

Answer by ewatrrr(24785)

 
Hi,
password containing six numerals to be able to use the site:
no digit may be used more than once. How many passwords are allowed
10*9*8*7*6*5 = 151,200
Note: 10P6 is the same as the above(using permutation as order is important)
10P6 = 10!/(10-6)! = 10!/4! = 10*9*8*7*6*5